OpenStudy (anonymous):

simple interest formula? 30000 earned 500 in 2years. what is the interest rate? sd

OpenStudy (anonymous):

PERT

OpenStudy (anonymous):

y = P*e^(r*t)

OpenStudy (anonymous):

still don't know what e^ is.

OpenStudy (anonymous):

step 1. 500 = 30000*e^(r*2) step 2. 0.0166666667 = e^(2r) step 3. ln (0.01667) = ln (e^(2r)) step 4. ln (0.01667) = 2r step 5. r = [ln(0.01667)]/2

OpenStudy (anonymous):

e is exponential ln is natural log

OpenStudy (anonymous):

that's not simple interest, that's continuous compounding interest

OpenStudy (anonymous):

500 = 30000 * 2 * r r = 500 / 60000 = .083333 = 8.3333..%

OpenStudy (anonymous):

thx . that's what i was looking for!

OpenStudy (anonymous):

except i am getting .008 or.8%?